OpenStudy (shaleiah):

A right triangle has leg lengths 0.3 centimeters and 0.4 centimeters. What is cosθ where θ is the smallest angle of the triangle?

OpenStudy (aravindg):

Cos theta equals adjacent side divided by hypotenuse

OpenStudy (shaleiah):

So, \[\cos ^{-1}(\frac{ 0.3 }{ 0.4 })\] ?

OpenStudy (aravindg):

You need to first calculate hypotenuse using pythagorean theorem

OpenStudy (shaleiah):

a^2+b^2=c^2 0.3^2+0.4^=c^2 sqrt 0.25 0.5

OpenStudy (shaleiah):

@AravindG

OpenStudy (aravindg):

Good.

OpenStudy (aravindg):

Now for smallest angle adjacent side will be smallest of .4 and .3 and u gt the hypotenuse. cant u find cos theta now?

OpenStudy (shaleiah):

0.7227342478

OpenStudy (aravindg):

Good.